OpenStudy (anonymous):

Need help with the washer method: x=1-y^2, x=2+y^2, =-1, y=1

OpenStudy (anonymous):

*should be y=-1

OpenStudy (anonymous):

x=-1 , y=-1?

OpenStudy (anonymous):

try out this website. http://tutorial.math.lamar.edu/Classes/CalcII/CalcII.aspx

OpenStudy (anonymous):

So you have two function. Find out which one is to the right.

OpenStudy (anonymous):

x=2+y^2 would be to the right, right?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

are your functions correct, they don't seem to intersect

OpenStudy (anonymous):

yep, they are correct. I also forgot to put its revolved around the y-axis

OpenStudy (anonymous):

because the functions don't even touch each other

OpenStudy (anonymous):

That's really strange, I wonder if the text made an error

OpenStudy (anonymous):

For washer method, we are supposed to take revolution of intersection between two functions

OpenStudy (anonymous):

That much I understand, maybe I should try converting the equation to xs and see what happens.

OpenStudy (anonymous):

If we change second function to x=-2+y^2 we can solve it, do you wanna

OpenStudy (anonymous):

sure : )

OpenStudy (anonymous):

We will rotate around x=-3

OpenStudy (anonymous):

Here come the integral \[\pi \int\limits_{?}^{?}((-y^2-1)-(y^2-2)+1)^2-1\]

OpenStudy (anonymous):

dy

OpenStudy (anonymous):

did you get it, mathrocks?

OpenStudy (anonymous):

Ok, this is starting to make some sense.

OpenStudy (anonymous):

you have to find points of intersection though

OpenStudy (anonymous):

Never memorize any formula for disc,shell, washer method

OpenStudy (anonymous):

I can find the intersection points with no problem. It's just setting up the washer equation that sometimes trips me up. Thanks for your help, I appreciate it.

OpenStudy (anonymous):

just know that when you are doing this type of problem, you are adding up circle

OpenStudy (anonymous):

If I helped you in somehow, can you fan me

OpenStudy (anonymous):

I'll remember that and will practice drawing the actual graphs. And the fan request has been done. : )